(iii) During a laboratory experiment, the average number of radioactive particles passing through   a counter in 1 millisecond is 2. What is the probability that 5 particles enter the counter in any given millisecond?                                                                                                       



1
Expert's answer
2021-07-16T11:57:16-0400

Poisson distribution

"P(X=x) = \frac{e^{-λ} \times λ^x }{x!} \\

λ=2 \\

x=5 \\

P(X=5) = \frac{e^{-2} \times 2^5 }{5!} \\

= 0.03609"